2. (25 pts) Suppose you find two stocks in the market that they are perfectly negatively correlated (p. They have the following characteristics: Mean E(r) Stock x0.10 Stock y 0.16 St.Dev, σ 0.05 0.105 (a) (10 pts) If you want to form a portfolio by only the above stocks, what would be the proportion (weight) of your money invested in each stock so you can achieve the lowest possible risk (minimum variance portfolio)? (b) (10 pts) What is the expected return and the standard deviation of the minimumvar ance portfolio from (a)?
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text ta copy. B า 프 . Ej-., Δ. : r_一 逻锂函Merge & Center. $, % , 弼,8 conditional Format . Cell Insert Delete Format Paste Sort &Find & Format Painter Formatting as Table Styles2 Clear Clipboard BG57 BF Font Alignment Number Cells Edting Formula Bar BG BH BI BJ BK BL BN BO BP BQ BR BS BT BU 42 43 45 46 47 48 49 50 51 52 53 54 σ (A) σ (B) (o (A) )A2 25 (o (B))A2 110.25 1.0000 5.0000 TAKE A = X 10.5000 TAKE B-Y W(A)(110.25) ((1F5)10.5 (25) (110.25) (2*(-1)*(5) (10.5)) 56 57 58 59 60 i1 1 240.25 WEIGHTS WIA- w(x) w(B)= w(Y) 0.6774 67.74% 0.3226 32.26% İ İ MIN VAR PORT C BEST CAL BİNOMIAL 2 TIME -SİNKING AMORTIZA A Sheet2 JORDAN INSURANCE 14:18 31-01-2019Home nert Page Layout Formulas Data Review View dd-Ins s Cut aCopy Σ AutoSum ー E ゴWrap Text в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 Conditional Format eCell Insert Delete Format Paste Sort &Find & Format Painter Formatting as Table Styles2 Clear Clipboard BG71 BG Font Alignment Number Styles Cells Edting BH BI BJ BK BL BM BN BO BP BQ BR BS BT 57 58 a 59 60 61 62 63 64 65 WEIGHTS W(A )= w(X) 0.6774 67.74% 0.3226 32.26% W(B)W(Y) W(A) R(A) = PORTFOLIO RETURN W(A)*R(A) W(B)*R(B) PORTFOLIO RETURN = B161*B162 + BM61*SM62 = 67.74% 10% w(B)-32.26% A(B) = 16% 11.94% PORTFOLIO VARIANCE (W(A))A2 (SD(A))A2 (W(B))A2*(SD(B))A2 2 W(A) W(B) SD(A) SD (B) r PORTFOLIO VARIANCE= [(0.6774)2(5)^2 +(0.3226)2(10.5)^2 +2*(0.6774)(0.3226)(5)*(10.5)(-1)] PORTFOLIO VARIANCE- PORTFOLIO STANDARD DEVIATION = 67 68 69 70 71 72 73 74 75 il 1 İ İ MIN VAR PORT C BEST CAL BİNOMIAL 2 TIME --SİNKING AMORTIZA . Sheet2 0.000 % 0.000 % JORDAN INSURANCE 31-01-2019